OpenStudy (anonymous):

find the sum. 7+10+13+...+97

OpenStudy (anonymous):

The sequence 7,10,13,...97 is a arithmetic sequence, with \[\Large a_1=7\\ \Large d=3\] And if n is the number of the term 97 then : \[\Large 97= a_n=a_1+(n-1)d=7+3(n-1)\] So : \[\Large 3(n-1)=97-7=90\] So : \[\Large n-1=30\] So : \[\Large n=31\] Now to find the sum we have : \[\Large 7+10+13+\cdots+97=a_1+a_2+\cdots+a_{31} \\\Large~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=\frac{31}{2}(a_1+a_{31}) \\\Large~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=\frac{31}{2}(3+97) \\\Large~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~=1550 \]